### v1

# Computing Richelot isogeny graph of superspecial abelian threefolds

Abstract.
In algebraic geometry, superspecial curves are important research objects. While the number of superspecial genus-3 curves in characteristic $p$ is known, the number of hyperelliptic ones among them has not been determined even for small $p$. In this paper, in order to compute the latter number, we give an algorithm for computing the $(2,2,2)$-isogeny graph of superspecial abelian threefolds by using theta functions. This enables us to list up all superspecial genus-3 curves efficiently. Executing this algorithm on Magma, we succeeded in counting all superspecial hyperelliptic curves of genus $3$ when $10011\leq p<$ 100, and in finding a single one when $100007\leq p<$ 10000.
1. Introduction
Throughout this paper, by a curve we mean a non-singular projective variety of dimension one and isomorphisms between two curves are ones over an algebraically closed field. A curve $C$ over an algebraically closed field $k$ of characteristic $p>0$ is said to be superspecial if and only if the Jacobian of $C$ is isomorphic to the product of supersingular elliptic curves. Recently, superspecial curves are used for isogeny-based cryptography (e.g. [16], [5], [1] and so on), and then studies on such curves are becoming more important.
For given integers $g$ and $p$, it is known that the number of isomorphism classes of superspecial curves of genus $g$ in characteristic $p$ is finite, and determining that number is an important problem. This problem was already solved for genus $g\leq 3$; Deuring [11] found the number of all supersingular elliptic curves (i.e. $g=1$), and the number for $g=2$ is given by Ibukiyama, Katsura, and Oort [21]. For the case of $g=3$, Oort [34] and Ibukiyama [20] showed there exists a superspecial curve of genus $3$ in arbitrary characteristic $p\geq 3$, and the number of such curves was also computed by Brock [3]. However, the number of a superspecial hyperelliptic curve of genus 3 is not known, and in fact, even the existence of such a curve in arbitrary characteristic $p\geq 3$ has not been shown.
[…]
Consequently, we obtain the following main theorem:
Theorem 1.1.
Under some assumptions (see Remark 3.11 and Assumption 3.13), there is an algorithm (Algorithm 3.7) for computing the graph $\mathcal{G}_{3}(2,p)$ with $\widetilde{O}(p^{6})$ operations in the finite field $\hskip 0.85358pt\mathbb{F}_{p^{2}}$ for all chracteristics $p>7$.
Our algorithm allows us to restore the explicit defining equations of all superspecial curves of genus 3 efficiently, compared to the methods using the Gröbner basis (this complexity would be exponential with respect to $p$ in worst case).
Executing our algorithm, we succeeded in solving the existence of superspecial hyperelliptic curves of genus 3 in small $p$ as follows:
Theorem 1.2.
There exists a superspecial hyperelliptic curve of genus 3 in characteristic $p$ with $100007\leq p<$ 10000. In addition, the numbers of isomorphism classes of such curves for $10011\leq p<$ 100 are determined (as in Table 1 of Section 4.1).
We note that the upper bounds on $p$ in Theorem 1.2 can be increased easily. From our computational results, we obtain a conjecture that there exists a superspecial hyperelliptic curve of genus 3 in arbitrary characteristic $p\geq 7$.
The rest of this paper is organized as follows:  Section 2 is dedicated to reviewing some facts on analytic theta functions. In Section 2.4, we recall the definition and properties of superspecial Richelot isogeny graph $\mathcal{G}_{g}(2,p)$. We then explain how to compute the graph $\mathcal{G}_{3}(2,p)$ in Section 3, and finally Algorithm 3.7 gives its explicit construction. In Section 4, we state computational results about the existence and the number of superspecial hyperelliptic curves of genus 3 in small $p$, obtained by our implementation. At last, we give a concluding remark in Section 5.
Acknowledgements
This research was conducted under a contract of “Research and development on new generation cryptography for secure wireless communication services” among “Research and Development for Expansion of Radio Wave Resources (JPJ000254)”, which was supported by the Ministry of Internal Affairs and Communications, Japan. This work was supported by JSPS Grant-in-Aid for Young Scientists 20K14301 and 23K12949, and WISE program (MEXT) at Kyushu University.
2. Preliminaries
In this section, we summarize background knowledge required in later sections. Abelian varieties and isogenies between them until Section 2.3 are defined over $\mathbb{C}$, but in Section 2.4 we consider those over a field of odd characteristic.
2.1. Abelian varieties and their isogenies
In this subsection, we review about complex abelian varieties and isogenies. A complex abelian variety  of dimension $g$ is isomorphic to $\mathbb{C}^{g}/(\mathbb{Z}^{g}+\varOmega\mathbb{Z}^{g})$ where $\varOmega$ is an element of the Siegel upper half-space
$\mathcal{H}_{g}\coloneqq\{\varOmega\in{\rm Mat}_{g}(\mathbb{C})\,\mid\,^{t}\varOmega=\varOmega,\,{\rm Im}\,\varOmega>0\}.\vspace{0.5mm}$
Then, we call $\varOmega\in\mathcal{H}_{g}$ a period matrix of the abelian variety. First of all, we recall a characterization of isomorphisms between abelian varieties:
Definition 2.1.
For a commutative ring $R$ and an integer $g\geq 1$, we define a group
${\rm Sp}_{2g}(R)\coloneqq\{M\in{\rm GL}_{2g}(R)\,\mid\,ME\hskip 0.85358pt^{t}M=E\},\quad E\coloneqq\begin{pmatrix}0&{\mathbf{1}}_{g}\\ -{\mathbf{1}}_{g}&0\end{pmatrix}.$
We say that a matrix $M\in{\rm GL}_{2g}(R)$ is symplectic over $R\hskip 0.56905pt$ if $M\in{\rm Sp}_{2g}(R)$.
The following lemma helps determine whether $M$ is a symplectic matrix or not:
Lemma 2.2 ([28, Lemma 8.2.1]).
For any $(2g\times 2g)$-matrix $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm GL}_{2g}(R)$ with $\alpha,\beta,\gamma,\delta\in{\rm Mat}_{g}(R)$, the following three statements are equivalent:
  1. (1)
    The matrix $M$ is symplectic over $R$.
  2. (2)
    Two matrices $\alpha\hskip 0.56905pt^{t}\beta,\,\gamma\hskip 0.85358pt^{t}\delta$ are both symmetric and $\alpha\hskip 0.85358pt^{t}\delta-\beta\hskip 1.42262pt^{t}\gamma={\mathbf{1}}_{g}$.
  3. (3)
    Two matrices ${}^{t}\alpha\hskip 0.85358pt\gamma,\hskip 1.42262pt^{t}\beta\hskip 1.42262pt\delta$ are both symmetric and ${}^{t}\alpha\hskip 0.85358pt\delta-\hskip 1.42262pt^{t}\beta\hskip 1.42262pt\gamma={\mathbf{1}}_{g}$.
For a matrix $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm Sp}_{2g}(\mathbb{Z})$, the action on the Siegel upper half-space
$\begin{split}{\rm Sp}_{2g}(\mathbb{Z})\times\mathcal{H}_{g}&\longrightarrow\mathcal{H}_{g}\\[-1.42262pt] (M,\varOmega)&\longmapsto(\alpha\varOmega+\beta)(\gamma\varOmega+\delta)^{-1}\eqqcolon M.\varOmega\end{split}$
is well-defined. Moreover, the map
$\mathbb{C}^{g}\longrightarrow\mathbb{C}^{g}\,;\,z\longmapsto^{t}(\gamma\varOmega+\delta)^{-1}z\eqqcolon M.z\vspace{0.5mm}$
induces an isomorphism $\mathbb{C}^{g}/(\mathbb{Z}^{g}+\varOmega\mathbb{Z}^{g})\rightarrow\mathbb{C}^{g}/(\mathbb{Z}^{g}+M.\varOmega\mathbb{Z}^{g})$.
Theorem 2.3 ([28, Theorem 8.3.1]).
Two matrices $\varOmega,\varOmega^{\prime}\in\mathcal{H}_{g}$ determine isomorphic abelian varieties if and only if there exists $M\in{\rm Sp}_{2g}(\mathbb{Z})$ such that $\varOmega^{\prime}=M.\varOmega$.
An isogeny is a surjective homomorphism between two abelian varieties. If there exists an isogeny $A\rightarrow B$, then the dimensions of $A$ and $B$ are equal to each other, and we say that $A$ and $B$ are isogenous. Two isogenies $A\rightarrow B$ with the same kernel are equivalent up to an automorphism of $B$, and thus we identify them.
Proposition 2.4.
For an abelian variety $A$ of dimension $g$ and a prime integer $\ell$, the number of maximal isotropic subgroups $G$ of $A[\ell]$ is equal to
$N_{g}(\ell)\coloneqq\prod_{k=1}^{g}(\ell^{k}+1).$
Proof.
See [8, Lemma 2] or [4, Lemma 2]. ∎
An isogeny whose kernel is such a group $G$ the above is called an $(\overbrace{\ell,\dots,\ell}^{g})$-isogeny. In particular, one can see that the $(\ell,\ldots,\ell)$-isogeny
$\begin{split}\mathbb{C}^{g}/(\mathbb{Z}^{g}+\varOmega\mathbb{Z}^{g})&\longrightarrow\mathbb{C}^{g}/(\mathbb{Z}^{g}+\ell\varOmega\mathbb{Z}^{g})\\[-1.42262pt] z&\longmapsto\ell z\end{split}$
has the kernel $\frac{1}{\ell}\mathbb{Z}^{g}/\mathbb{Z}^{g}$.
2.2. Theta functions
The theta function of characteristics $a,b\in\mathbb{Q}^{g}$ is given as
[…]
for $z\in\mathbb{C}^{g}$ and $\varOmega\in\mathcal{H}_{g}$. All the characteristics can be considered modulo $\mathbb{Z}^{g}$ since one can check that
$\theta\bigl{[}\begin{smallmatrix}a+\alpha\\ b+\beta\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)\exp(2\pi i\hskip 1.42262pt^{t}a\beta)$
for all $\alpha,\beta\in\mathbb{Z}^{g}$. Then, in the paper, we consider only theta functions $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)$ for characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$.
Lemma 2.5.
For all characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$, the function $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)$ is even with respect to $z$ if and only if $4\hskip 0.56905pt^{t}ab$ is an even integer;  otherwise $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)$ is odd.
Proof.
The first assertion follows from that
$\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(-z,\varOmega)=(-1)^{4\hskip 0.56905pt^{t}ab}\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega).$
The second assertion follows from that $4\hskip 0.56905pt^{t}ab$ is an integer for all $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$. ∎
The following two formulas are very fundamental and generate many relations:
Theorem 2.6 (Riemann’s theta formula).
Let $m_{k}:=\begin{pmatrix}a_{k}\\ b_{k}\end{pmatrix}$ with $a_{k},b_{k}\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$, and we define
$\begin{pmatrix}n_{1}\\ n_{2}\\ n_{3}\\ n_{4}\end{pmatrix}\coloneqq\hskip 0.85358pt\frac{1}{2}\begin{pmatrix}{\mathbf{1}}_{2g}&\hskip 7.68222pt{\mathbf{1}}_{2g}&\hskip 7.68222pt{\mathbf{1}}_{2g}&\hskip 7.68222pt{\mathbf{1}}_{2g}\\ {\mathbf{1}}_{2g}&\hskip 7.68222pt{\mathbf{1}}_{2g}&-{\mathbf{1}}_{2g}&-{\mathbf{1}}_{2g}\\ {\mathbf{1}}_{2g}&-{\mathbf{1}}_{2g}&\hskip 7.68222pt{\mathbf{1}}_{2g}&-{\mathbf{1}}_{2g}\\ {\mathbf{1}}_{2g}&-{\mathbf{1}}_{2g}&-{\mathbf{1}}_{2g}&\hskip 7.68222pt{\mathbf{1}}_{2g}\end{pmatrix}\begin{pmatrix}m_{1}\\ m_{2}\\ m_{3}\\ m_{4}\end{pmatrix}.$
Then, we have the equation
$\prod_{k=1}^{4}\theta[m_{k}](0,\varOmega)=\frac{1}{2^{g}}\sum_{\alpha,\beta\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\prod_{k=1}^{4}\theta[n_{k}+\gamma_{\alpha,\beta}](0,\varOmega)\vspace{0.5mm}$
where $\gamma_{\alpha,\beta}$ denotes $\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$ with $\alpha,\beta\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$.
Proof.
Apply [23, Chapter IV, Theorem 1] for $z_{1}=z_{2}=z_{3}=z_{4}=0$. ∎
Theorem 2.7 (Duplication formula).
For all $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$, we have
$\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega)^{2}=\frac{1}{2^{g}}\hskip 0.56905pt\sum_{\beta\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}0\\ b\end{smallmatrix}\bigr{]}(z,\varOmega/2)\theta\bigl{[}\begin{smallmatrix}0\\ b+\beta\end{smallmatrix}\bigr{]}(z,\varOmega/2).$
Proof.
Apply [23, Chapter IV, Theorem 2] for $m_{1}=m_{2}$ and $z_{1}=z_{2}=z$. ∎
Especially if a matrix $\varOmega\in\mathcal{H}_{g}$ is written as
$\varOmega=\begin{pmatrix}\varOmega_{1}&0\\ 0&\varOmega_{2}\end{pmatrix}\ \,{\rm where}\ \,\varOmega_{1}\in\mathcal{H}_{g_{1}},\ \varOmega_{2}\in\mathcal{H}_{g_{2}}$
with $g_{1}+g_{2}=g$, then we have the following lemma:
Lemma 2.8 ([9, Lemma F.3.1]).
With notations the above, we write $z=(z_{1},z_{2})$ where $z_{1}\in\mathbb{C}^{g_{1}},\,z_{2}\in\mathbb{C}^{g_{2}}$. Then, we have the equation
$\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}\\ b_{1}&b_{2}\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a_{1}\\ b_{1}\end{smallmatrix}\bigr{]}(z_{1},\varOmega_{1})\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}a_{2}\\ b_{2}\end{smallmatrix}\bigr{]}(z_{2},\varOmega_{2})\vspace{0.5mm}$
for all characteristics $a_{1},b_{1}\in\frac{1}{2}\mathbb{Z}^{g_{1}}/\mathbb{Z}^{g_{1}}$ and $a_{2},b_{2}\in\frac{1}{2}\mathbb{Z}^{g_{2}}/\mathbb{Z}^{g_{2}}$.
Let us return to cases with a general (not necessarily diagonal) matrix $\varOmega\in\mathcal{H}_{g}$. For any symplectic matrix $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm Sp}_{2g}(\mathbb{Z})$ and characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$, we define a scalar
[…]
and two vectors
(2.1) $M.a:=a\hskip 0.85358pt^{t}\delta-b\hskip 1.13809pt^{t}\gamma+\frac{1}{2}(\gamma\hskip 1.42262pt^{t}\delta)_{0},\quad M.b:=b\hskip 1.42262pt^{t}\alpha-a\hskip 0.85358pt^{t}\beta+\frac{1}{2}(\alpha\hskip 0.85358pt^{t}\beta)_{0},\vspace{0.5mm}$
where $(\alpha\hskip 0.85358pt^{t}\beta)_{0},(\gamma\hskip 1.42262pt^{t}\delta)_{0}$ denote the row vector of the diagonal elements of $(\alpha\hskip 0.85358pt^{t}\beta),(\gamma\hskip 1.42262pt^{t}\delta)$. Then, the matrix $M$ acts on the theta function in the following way:
Theorem 2.9 (Theta transformation formula).
For all $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$, we have
$\theta\bigl{[}\begin{smallmatrix}M.a\\ M.b\end{smallmatrix}\bigr{]}(M.z,M.\varOmega)=c\cdot e^{\pi ik(M,a,b)}\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(z,\varOmega),\vspace{1mm}$
where $c\in\mathbb{C}^{\times}$ is a constant which does not depend $a$ or $b$.
Proof.
This follows from [28, Theorem 8.6.1]. ∎
Let $n\geq 1$ be an integer and we define
$\varGamma_{n}$ :=$:=\bigl{\{}M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm Sp}_{2g}(\mathbb{Z})\,\mid\,\alpha\equiv\delta\equiv{\mathbf{1}}_{g},\ \beta\equiv\gamma\equiv{\mathbf{0}}_{g}\pmod{n}\bigr{\}},$
$\varGamma_{n,2n}$ :=$:=\bigl{\{}M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in\varGamma_{n}\,\mid\,{\rm diag}(\hskip 0.85358pt^{t}\alpha\gamma)\equiv{\rm diag}(\hskip 0.85358pt^{t}\beta\hskip 0.85358pt\delta)\equiv 0\pmod{2n}\},$
which are both subgroups of ${\rm Sp}_{2g}(\mathbb{Z})$.
Corollary 2.10.
For all characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{g}/\mathbb{Z}^{g}$, we have
$\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,M.\varOmega)^{4}}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,M.\varOmega)^{4}}$ $=\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)^{4}}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\varOmega)^{4}}\quad{\rm if}\ M\in\varGamma_{2},$
$\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,M.\varOmega)^{2}}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,M.\varOmega)^{2}}$ $=\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)^{2}}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\varOmega)^{2}}\quad{\rm if}\ M\in\varGamma_{2,4},$
$\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,M.\varOmega)}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,M.\varOmega)}$ $=\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\varOmega)}\quad\ \,{\rm if}\ M\in\varGamma_{4,8}.$
Proof.
It follows from Theorem 2.9 that
$\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,M.\varOmega)}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,M.\varOmega)}=\frac{\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\varOmega)}\exp\bigl{(}\pi ia\hskip 0.85358pt^{t}\beta\hskip 0.85358pt\delta\hskip 1.42262pt^{t}a-\pi ib\hskip 0.85358pt^{t}\gamma\hskip 0.85358pt\alpha\hskip 1.42262pt^{t}b-2\pi ia(\hskip 0.56905pt^{t}\alpha-\mathbf{1}_{g})\hskip 0.56905pt^{t}b\bigr{)}.$
Hence, we obtain this assertion by simple computations. ∎
2.3. Theta functions for abelian threefolds
In this subsection, we focus only on the case that $A$ is an abelian threefold (i.e.  abelian variety of dimension $g=3$), and we discuss theta functions associated to $A$. According to [28, Corollary 11.8.2], such $A$ is either of the following 4 types.
$\begin{split}\left\{\begin{array}[]{l}{\rm Type\ 1\!:\ }\,A\cong{\rm Jac}(C)\ \text{where}\ C\ \text{is\ a\ smooth\ plane\ quartic},\\[1.42262pt] {\rm Type\ 2\!:\ }\,A\cong{\rm Jac}(C)\ \text{where}\ C\ \text{is\ a\ hyperelliptic\ curve\ of\ genus\ }3,\\[1.42262pt] {\rm Type\ 3\!:\ }\,A\cong E\times{\rm Jac}(C)\ \text{where}\ E\ \text{and}\ C\text{\ is\ a genus-1 and genus-2 curve},\\[1.42262pt] {\rm Type\ 4\!:\ }\,A\cong E_{1}\times E_{2}\times E_{3}\ \text{where}\ E_{1},E_{2}\text{\ and\ }E_{3}\text{\ are\ elliptic\ curves}.\end{array}\right.\vspace{0.5mm}\end{split}$
First of all, we introduce the following notations for the sake of simplicity.
Definition 2.11.
For a period matrix $\varOmega\in\mathcal{H}_{3}$ of $A$, we write
$\vartheta_{i}(z,\varOmega):=\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z,\varOmega),\quad i:=2b_{1}+4b_{2}+8b_{3}+16a_{1}+32a_{2}+64a_{3},\vspace{0.5mm}$
where $a_{1},a_{2},a_{3},b_{1},b_{2}$, and $b_{3}$ belong to $\{0,1/2\}$. Note that $i$ runs over $\{0,\ldots,63\}$. Moreover, the value $\vartheta_{i}(0,\varOmega)$ is simply be denoted as $\vartheta_{i}$.
We consider an abelian threefold $A\cong\mathbb{C}^{3}/(\mathbb{Z}^{3}+\varOmega\mathbb{Z}^{3})$, then the projective value
$[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}(\mathbb{C})\vspace{0.5mm}$
is called a theta null-point of $A$. Remark that there are different theta null-points for isomorphic abelian threefolds.
Definition 2.12.
We define the sets
[…]
and $I_{\rm odd}\coloneqq\{0,\ldots,63\}\smallsetminus I_{\rm even}$. Here, we remark that #$\#I_{\rm even}=36$ and #$\#I_{\rm odd}=28$.
It follows from Lemma 2.5 that $\vartheta_{i}(z)$ is even (resp. odd) if and only if $i\in I_{\rm even}$ (resp. $i\in I_{\rm odd}$). Therefore, we have that $\vartheta_{i}=0$ for all $i\in I_{\rm odd}$.
Definition 2.13.
Let $S_{\rm van}$ be the set of vanishing indices $i\in I_{\rm even}$, that is,
$S_{\rm van}\coloneqq\{\hskip 0.56905pti\in I_{\rm even}\mid\vartheta_{i}=0\}.$
In addition, we denote by $N_{\rm van}$ the cardinality of the set $S_{\rm van}$.
Lemma 2.14.
The value $N_{\rm van}$ is invariant for isomorphic abelian threefolds.
Proof.
This follows from the theta transformation formula (Theorem 2.9). ∎
Then, we can identify by $N_{\rm van}$ which of the above 4 types $A$ corresponds (remark that this does not depend on a period matrix $\varOmega$ of $A$ by the above lemma).
Proposition 2.15.
The possible values of $N_{\rm van}$ are $0,1,6,9$ and moreover
[…]
Proof.
The first assertion follows from [18, Theorem 3.1] (the cases where $N_{\rm van}>9$ correspond to singular abelian threefolds). The second assertion for $N_{\rm van}=0,1$ are shown by [35, Theorem 2], and hence we show the remaining cases.
Let $A$ be an abelian threefold such that $A$ is isomorphic to the product $A_{1}\times A_{2}$ of an elliptic curve $A_{1}$ and an abelian surface $A_{2}$. Denote $\varOmega_{1},\varOmega_{2}$ and $\varOmega$ be one of period matrices of $A_{1},A_{2}$ and $A$, respectively. Then, there exists $M\in{\rm Sp}_{6}(\mathbb{Z})$ such that $M.\varOmega={\rm diag}(\varOmega_{1},\varOmega_{2})$ from Theorem 2.3, but we may assume $\varOmega={\rm diag}(\varOmega_{1},\varOmega_{2})$ by using Lemma 2.14. It follows from Lemma 2.8 that
$\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z,\varOmega)=\theta\bigl{[}\begin{smallmatrix}a_{1}\\ b_{1}\end{smallmatrix}\bigr{]}(z_{1},\varOmega_{1})\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}a_{2}&a_{3}\\ b_{2}&b_{3}\end{smallmatrix}\bigr{]}(z_{2},\varOmega_{2}),\quad z\coloneqq(z_{1},z_{2}),\vspace{0.5mm}$
and hence we see that $N_{\rm van}=6$ (resp.  $9$) if $A_{2}$ is simple (resp. not simple). ∎
At the rest of this subsection, we assume that an abelian threefold $A$ corresponds to a genus-3 curve $C$. Let us review how to restore a defining equation of $C$ in the case that $C$ is hyperelliptic.
Proposition 2.16.
We assume that $S_{\rm van}=\{61\}$ for a period matrix $\varOmega\in\mathcal{H}_{3}$ of $A$. Then $A$ is isomorphic to the Jacobian of a genus-3 hyperelliptic curve
$C:y^{2}=x(x-1)(x-\lambda_{1})(x-\lambda_{2})(x-\lambda_{3})(x-\lambda_{4})(x-\lambda_{5}),$
where
(2.2) $\lambda_{1}\coloneqq\frac{\vartheta_{42}^{2}\vartheta_{2}^{2}}{\vartheta_{5}^{2}\vartheta_{45}^{2}},\ \,\lambda_{2}\coloneqq\frac{\vartheta_{42}^{2}\vartheta_{3}^{2}}{\vartheta_{4}^{2}\vartheta_{45}^{2}},\ \,\lambda_{3}\coloneqq\frac{\vartheta_{27}^{2}\vartheta_{42}^{2}}{\vartheta_{45}^{2}\vartheta_{28}^{2}},\ \,\lambda_{4}\coloneqq\frac{\vartheta_{2}^{2}\vartheta_{49}^{2}}{\vartheta_{54}^{2}\vartheta_{5}^{2}},\ \,\lambda_{5}\coloneqq\frac{\vartheta_{3}^{2}\vartheta_{0}^{2}}{\vartheta_{7}^{2}\vartheta_{4}^{2}}.$
Proof.
Apply Thomae’s formula (cf. [39]) for $g=3$ or see [36, Theorem 4]. ∎
On the other hand, even in the case that $C$ is a plane quartic, we can also restore an explicit equation of $C$. Taking a period matrix $\varOmega\in\mathcal{H}_{3}$ of $A={\rm Jac}(C)$, then it follows from Proposition 2.15 that $\vartheta_{i}\neq 0$ for all $i\in I_{\rm even}$. We define the following non-zero values
$a_{11}$ $\coloneqq\frac{\vartheta_{12}\vartheta_{5}}{\vartheta_{40}\vartheta_{33}},$ $\quad a_{21}$ $\coloneqq\frac{\vartheta_{27}\vartheta_{5}}{\vartheta_{40}\vartheta_{54}},$ $\quad a_{31}$ $\coloneqq-\frac{\vartheta_{27}\vartheta_{12}}{\vartheta_{33}\vartheta_{54}},$
(2.3) $a_{12}$ $\coloneqq\frac{\vartheta_{21}\vartheta_{28}}{\vartheta_{49}\vartheta_{56}},$ $\quad a_{22}$ $\coloneqq\frac{\vartheta_{2}\vartheta_{28}}{\vartheta_{49}\vartheta_{47}},$ $\quad a_{32}$ $\coloneqq\frac{\vartheta_{2}\vartheta_{21}}{\vartheta_{56}\vartheta_{47}},$
$a_{13}$ $\coloneqq\frac{\vartheta_{7}\vartheta_{14}}{\vartheta_{35}\vartheta_{42}},$ $\quad a_{23}$ $\coloneqq\frac{\vartheta_{16}\vartheta_{14}}{\vartheta_{35}\vartheta_{61}},$ $\quad a_{33}$ $\coloneqq\frac{\vartheta_{16}\vartheta_{7}}{\vartheta_{42}\vartheta_{61}}$
computed as in [13, p. 15]. Then, there exist $k_{1},k_{2},k_{3},\lambda_{1},\lambda_{2},\lambda_{3}\in\mathbb{C}^{\times}$ such that
(2.4) $\begin{pmatrix}\lambda_{1}a_{11}&\lambda_{2}a_{21}&\lambda_{3}a_{31}\\[1.42262pt] \lambda_{1}a_{12}&\lambda_{2}a_{22}&\lambda_{3}a_{32}\\[1.42262pt] \lambda_{1}a_{13}&\lambda_{2}a_{23}&\lambda_{3}a_{33}\end{pmatrix}\begin{pmatrix}k_{1}\\[1.42262pt] k_{2}\\[1.42262pt] k_{3}\end{pmatrix}=\begin{pmatrix}1/a_{11}&1/a_{21}&1/a_{31}\\[1.42262pt] 1/a_{12}&1/a_{22}&1/a_{32}\\[1.42262pt] 1/a_{13}&1/a_{23}&1/a_{33}\end{pmatrix}\begin{pmatrix}\lambda_{1}\\[1.42262pt] \lambda_{2}\\[1.42262pt] \lambda_{3}\end{pmatrix}=\begin{pmatrix}-1\\[1.42262pt] -1\\[1.42262pt] -1\end{pmatrix}\!.$
Then, we can restore the curve $C$ from the values of $k_{i}$ using the following fact:
Theorem 2.17 (Weber’s formula).
In the above situation, there exist three homogeneous linear polynomials $\xi_{1},\xi_{2},\xi_{3}\in\mathbb{C}[x,y,z]$ such that
$\left\{\begin{array}[]{l}\xi_{1}+\xi_{2}+\xi_{3}+x+y+z=0,\\[1.42262pt] \displaystyle\frac{\xi_{1}}{a_{i1}}+\frac{\xi_{2}}{a_{i2}}+\frac{\xi_{3}}{a_{i3}}+k_{i}(a_{i1}x+a_{i2}y+a_{i3}z)=0,\quad{\rm for\ all}\ i\in\{1,2,3\}.\end{array}\right.$
Then $A$ is isomorphic to the Jacobian of a plane quartic
$C:(\xi_{1}x+\xi_{2}y-\xi_{3}z)^{2}-4\xi_{1}\xi_{2}xy=0.$
Proof.
[…]
2.4. Superspecial abelian threefold isogeny graph
Unlike the previous subsections, let us consider abelian varieties over a field of characteristic $p\geq 3$ in this subsection. An elliptic curve $E$ is called supersingular if $E[p]=\{0\}$.
Definition 2.18.
An abelian variety $A$ is called superspecial if $A$ is isomorphic to the product of supersingular elliptic curves (without a polarization). A superspecial curve is a curve whose Jacobian is a superspecial abelian variety.
We note that if two curves $C,C^{\prime}$ are not isomorphic to each other, then neither are their Jacobians (as principally polarized abelian varieties) from Torelli theorem. Let $\Lambda_{g,p}$ be the list of isomorphism classes of superspecial genus-$g$ curves over $\overline{\mathbb{F}}_{p}$. For $g\leq 3$, the number #$\#\Lambda_{g,p}$ is given by [11], [21], and [3], however it is not known how many of those curves are hyperelliptic, especially. In the following, we explain how to generate the list $\Lambda_{g,p}$ by using the isogeny graph. First of all, we define the isogeny graphs of superspecial abelian varieties:
Definition 2.19.
Let $\ell$ be a prime integer. The superspecial isogeny graph $\mathcal{G}_{g}(\ell,p)$ is the directed multigraph obtained as
  • The vertices are isomorphism classes of all principally polarized superspecial abelian varieties of dimension $g$ over $\overline{\mathbb{F}}_{p}$.
  • The edges are $(\ell,\ldots,\ell)$-isogenies defined over $\overline{\mathbb{F}}_{p}$ between two vertices.
Then, the graph $\mathcal{G}_{g}(\ell,p)$ is a $N_{g}(\ell)$-regular multigraph by Proposition 2.4.
In the following, we discuss the case where $\ell=2$, then the graph $\mathcal{G}_{g}(2,p)$ is often called the superspecial Richelot isogeny graph. The list $\Lambda_{g,p}$ can be generated using the superspecial Richelot isogeny graph as in Algorithm 2.1:
1:An integer $g\geq 2$ and a prime integer $p\geq 3$
2:The list of isomorphism classes of all superspecial curves of genus $g$
3:Generate the list $\Lambda_{1,p}$ of isom. classes of s.sing. elliptic curves
4:Let $\mathcal{L}$ be the list of isom. classes of $E_{1}\times\cdots\times E_{g}$ for all $E_{i}\in\Lambda_{1,p}$ and $k\leftarrow 1$
5:while $k\leq\#\mathcal{L}$ do
6:     Let $A$ be the $k$-th element of $\mathcal{L}$
7:     Compute all codomains of Richelot isogenies whose domain is $A$
8:     for all codomains $A^{\prime}$ obtained in the previous step do
9:         if $A^{\prime}$ is not isomorphic to any element of $\mathcal{L}$ then
10:              Add $A^{\prime}$ to the end of $\mathcal{L}$
11:         end if
12:     end for
13:     Let $k\leftarrow k+1$
14:end while
15:return the list of genus-$g$ curves $C$ such that the Jacobian of $C$ belongs to $\mathcal{L}$
Algorithm 2.1 ListUpSuperspecialCurves
The correctness of this algorithm can be checked thanks to the following theorem:
Theorem 2.20 ([24, Section 7]).
The graph $\mathcal{G}_{g}(\ell,p)$ is connected.
In the case of $g=2$, it is easy to implement Algorithm 2.1 (indeed, this is used in [25, Algorithm 5.1]). We can compute Richelot isogenies in Step 5 using formulas in [19, Proposition 4] and [38, Chapter 8]. Moreover, isomorphism tests in Step 7 can be done by computing Igusa invariants of a genus-2 curve (cf. [22]).
Remark 2.21.
The structure of the graph $\mathcal{G}_{2}(2,p)$ was studied by many authors (cf. [27], [15] and so on). An important fact is that the Jacobians’ subgraph $\mathcal{J}_{2}(2,p)$ of $\mathcal{G}_{2}(2,p)$ is also connected by [14, Theorem 7.2]. Here $\mathcal{J}_{g}(\ell,p)$ is defined to be
  • The vertices are isomorphism classes of all the Jacobians of genus-$g$ curves over $\overline{\mathbb{F}}_{p}$.
  • The edges are $(\ell,\ldots,\ell)$-isogenies defined over $\overline{\mathbb{F}}_{p}$ between two vertices.
These researches may make it possible for us to construct a more efficient algorithm for listing up all superspecial genus-2 curves rather than Algorithm 2.1.
Similarly to the case of $g=2$, we want to list up all genus-$g$ superspecial curves for $g\geq 3$. However, no explicit formula is known for the computations of $(2,\ldots,2)$-isogenies in general unlike the genus-2 case. In the next section, we will propose an algorithm for computing $(2,2,2)$-isogenies using theta functions.
3. Computing the Richelot isogeny graph of dimension $3$
In this section, we give an explicit algorithm (Algorithm 3.7) for computing the graph $\mathcal{G}_{3}(2,p)$. Our algorithm uses theta functions introduced in Section 2.2, but is still valid over a field of odd characteristic $p\geq 3$, thanks to the algebraic theory by Mumford [31]. The following three subsections describe the subroutines used in our main algorithm as follows:
  • Section 3.1: For a given theta null-point of an abelian threefold $A$, compute a theta null-point of one $A^{\prime}$, which is $(2,2,2)$-isogenous to $A$.
  • Section 3.2: For a given theta null-point of an abelian threefold $A$, compute theta null-points of all $A^{\prime}$, which are $(2,2,2)$-isogenous to $A$.
  • Section 3.3: For a given theta null-point of $A^{\prime}$ corresponding to a curve of genus 3, restore the defining equation of it.
At the last of this section (Section 3.4), we will combine them and write down our main algorithm explicitly.
3.1. How to compute one path from the vertex
In this subsection, let $A$ be a complex abelian threefold and $\varOmega\in\mathcal{H}_{3}$ be a period matrix of $A$. First of all, we prepare the following fact about a theta null-point of $A$.
Lemma 3.1.
For all matrices $\varOmega\in\mathcal{H}_{3}$, we have the equation
$\vartheta_{0}\vartheta_{1}\vartheta_{2}\vartheta_{3}-\vartheta_{4}\vartheta_{5}\vartheta_{6}\vartheta_{7}=\vartheta_{32}\vartheta_{33}\vartheta_{34}\vartheta_{35}$
with the notation in Definition 2.11.
Proof.
Applying Theorem 2.6 for four matrices
$m_{1}$ $\coloneqq\begin{pmatrix}0&0&0\\ 0&0&0\end{pmatrix},\quad$ $m_{2}$ $\coloneqq\begin{pmatrix}0&0&0\\ 1/2&0&0\end{pmatrix},$
$m_{3}$ $\coloneqq\begin{pmatrix}0&0&0\\ 0&1/2&0\end{pmatrix},\quad$ $m_{4}$ $\coloneqq\begin{pmatrix}0&0&0\\ 1/2&1/2&0\end{pmatrix},\vspace{0.5mm}$
then we obtain this lemma. ∎
We assume that a theta null-point $[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}(\mathbb{C})$ of $A$ is obtained. Squaring both sides of the equation in Lemma 3.1, we have
(3.5) $2\vartheta_{0}\vartheta_{1}\vartheta_{2}\vartheta_{3}\vartheta_{4}\vartheta_{5}\vartheta_{6}\vartheta_{7}=\vartheta_{0}^{2}\vartheta_{1}^{2}\vartheta_{2}^{2}\vartheta_{3}^{2}+\vartheta_{4}^{2}\vartheta_{5}^{2}\vartheta_{6}^{2}\vartheta_{7}^{2}-\vartheta_{32}^{2}\vartheta_{33}^{2}\vartheta_{34}^{2}\vartheta_{35}^{2},\vspace{0.5mm}$
which means that the value of $\prod_{j=0}^{7}\vartheta_{j}$ is determined by the given theta null-point. On the other hand, we can take any square root of ${\vartheta_{j}}^{2}$ for all $j\in\{0,\ldots,7\}$ so that the equation (3.5) holds, by the following proposition:
Proposition 3.2.
We suppose that $\vartheta_{0}\neq 0$ and let $k$ be an integer with $1\leq k\leq 6$. Then, there exists a symplectic matrix $M\in{\rm Sp}_{6}(\mathbb{Z})$ such that
[…]
for all $j\in\{1,\ldots,7\}$.
Proof.
It follows from Corollary 2.10 that all symplectic matrix in $\varGamma_{4,8}$ (resp. $\varGamma_{2,4}$) leave the values (resp. squares) of $\vartheta_{j}/\vartheta_{0}$ for all $j$. Let us construct $M\in\varGamma_{2,4}\!\smallsetminus\!\varGamma_{4,8}$ which changes the signs of only $\vartheta_{k}/\vartheta_{0}$ and $\vartheta_{7}/\vartheta_{0}$. We define
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(1,1,1,0,0,0)$ in the case of $k=1$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,1,0,1,1,0)$ in the case of $k=2$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,1,0,0,0,0)$ in the case of $k=3$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,0,1,0,1,1)$ in the case of $k=4$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,0,1,0,0,0)$ in the case of $k=5$,
  • $(c_{11},c_{12},c_{13},c_{22},c_{23},c_{33}):=(0,0,0,0,1,0)$ in the case of $k=6$
and $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}$ where
$\alpha\coloneqq\delta\coloneqq\mathbf{1}_{3},\quad\beta\coloneqq\mathbf{0}_{3},\quad\gamma\coloneqq\begin{pmatrix}4c_{11}&2c_{12}&2c_{13}\\ 2c_{12}&4c_{22}&2c_{23}\\ 2c_{13}&2c_{23}&4c_{33}\end{pmatrix}.\vspace{0.5mm}$
Then, one can check that $M$ belongs to ${\rm Sp}_{6}(\mathbb{Z})$ by Lemma 2.2, and moreover $M$ is the matrix we desire since
$\frac{\theta\bigl{[}\begin{smallmatrix}0\\ b\end{smallmatrix}\bigr{]}(0,M.\varOmega)}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,M.\varOmega)}=\frac{\theta\bigl{[}\begin{smallmatrix}0\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)}{\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\varOmega)}\exp(-\pi ib\gamma\hskip 1.42262pt^{t}b).\vspace{0.5mm}$
by Theorem 2.9. ∎
Summarizing the above discussions, we can compute $[\vartheta_{0}:\vartheta_{1}:\cdots:\vartheta_{7}]\in\mathbb{P}^{7}(\mathbb{C})$, from a given theta null-point of $A$. Note that if there exists an index $j\in\{0,\ldots,7\}$ such that ${\vartheta_{j}}^{2}=0$, then we can take any square root of $\vartheta_{k}^{2}$ for all $k\in\{0,\ldots,7\}\!\smallsetminus\!\{j\}$. Now, in the following, we consider the $(2,2,2)$-isogeny
$\begin{split}A=\mathbb{C}^{3}/(\mathbb{Z}^{3}+\varOmega\mathbb{Z}^{3})&\longrightarrow\mathbb{C}^{3}/(\mathbb{Z}^{3}+2\varOmega\mathbb{Z}^{3})=B\\[-1.42262pt] z&\longmapsto 2z\end{split}$
whose kernel is $\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}$, and we explain how to compute a theta null-point of the codomain $B$. Applying Theorem 2.7 for $z=0$ and $\varOmega\mapsto 2\varOmega$, we obtain
(3.6) $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,2\varOmega)^{2}=\frac{1}{2^{3}}\hskip 0.56905pt\sum_{\beta\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}}(-1)^{4\hskip 0.56905pt^{t}a\beta}\hskip 0.85358pt\theta\bigl{[}\begin{smallmatrix}0\\ b\end{smallmatrix}\bigr{]}(0,\varOmega)\theta\bigl{[}\begin{smallmatrix}0\\ b+\beta\end{smallmatrix}\bigr{]}(0,\varOmega)\vspace{0.5mm}$
for all characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}$. Since the right-hand side of (3.6) is obtained by the values $\vartheta_{0},\ldots,\vartheta_{7}$, we can compute a theta null-point of $B$. As a summary of the statements in this subsection, we give an explicit algorithm (Algorithm 3.2) to compute a null-point of the codomain of a $(2,2,2)$-isogeny:
1:A theta null-point $[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}$ of an abelian threefold $A$
2:A theta null-point of $B$, one of abelian threefolds $(2,2,2)$-isogenous to $A$
3:if any of $\vartheta_{0}^{2},\ldots,\vartheta_{7}^{2}$ is non-zero then
4:     Choose a square root of $(\vartheta_{k}/\vartheta_{0})^{2}$ for all $k\in\{1,\ldots,6\}$
5:     Compute $\vartheta_{7}/\vartheta_{0}$ from the equation (3.5)
6:else
7:     Find an index $j\in\{0,\ldots,7\}$ such that ${\vartheta_{j}}^{2}\neq 0$
8:     Choose a square root of $(\vartheta_{k}/\vartheta_{j})^{2}$ for all $k\in\{0,\ldots,7\}\!\smallsetminus\!\{j\}$
9:end if
10:for $i\hskip 0.85358pt\in\{0,\ldots,63\}$ do
11:     Write $i={a_{3}a_{2}a_{1}b_{3}b_{2}b_{1}}_{(2)}$ as a binary and initialize ${\vartheta^{\prime}}_{\!\!i}^{2}\leftarrow 0$
12:     for $j\in\{0,\ldots,7\}$ do
13:         Write $j={\beta_{3}\beta_{2}\beta_{1}}_{(2)}$ and set $k\leftarrow i+j\pmod{8}$
14:         Let ${\vartheta^{\prime}}_{\!\!i}^{2}\leftarrow{\vartheta^{\prime}}_{\!\!i}^{2}+(-1)^{a_{1}\beta_{1}+a_{2}\beta_{2}+a_{3}\beta_{3}}\cdot\vartheta_{j}\vartheta_{k}$
15:     end for
16:end for
17:return $[{\vartheta^{\prime}}_{\!\!0}^{2}:{\vartheta^{\prime}}_{\!\!1}^{2}:\cdots:{\vartheta^{\prime}}_{\!\!63}^{2}]\in\mathbb{P}^{63}$
Algorithm 3.2  ComputeOneRichelotIsogeny
We remark that there exists an index $j\in\{0,\ldots,7\}$ such that ${\vartheta_{j}}^{2}\neq 0$ in Step 5.
Indeed if $\vartheta_{0}=\vartheta_{1}=\cdots=\vartheta_{7}=0$, then it follows from (3.6) that $\theta\bigl{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\bigr{]}(0,2\varOmega)^{2}=0$ for all characteristics $a,b\in\frac{1}{2}\mathbb{Z}^{3}/\mathbb{Z}^{3}$, which contradicts Proposition 2.15.
Remark 3.3.
It suffices to do the loops from Step 8 to Step 14 only for $i\in I_{\rm even}$ since we have that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$ for all $i\in I_{\rm odd}$, in fact.
3.2. How to compute all paths from the vertex
We continue to assume that a theta null-point of a complex abelian threefold $A$ with the period matrix $\varOmega\in\mathcal{H}_{3}$ is given, and let $P,Q,R\in A$ as follows:
[…]
Then, the $(2,2,2)$-isogeny computed in Algorithm 3.2 has the kernel $G=\langle P,\hskip 0.85358ptQ,R\rangle$, which is determined by a given theta null-point of $A$. To compute different $(2,2,2)$-isogenies, we have to act an appropriate matrix $M\in{\rm Sp}_{6}(\mathbb{Z})$ on the theta null-point. Specifically, if we write $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}$ and let $M.P,M.Q,M.R\in A$ corresponding to
$(\alpha_{1j}/2,\alpha_{2j}/2,\alpha_{3j}/2)+\varOmega(\gamma_{1j}/2,\gamma_{2j}/2,\gamma_{3j}/2),\quad{\rm where}\ \alpha=(\alpha_{ij}),\ \gamma=(\gamma_{ij})\vspace{0.5mm}$
for $j\in\{1,2,3\}$ respectively, then we can compute the $(2,2,2)$-isogeny whose kernel is given as $M.G\coloneqq\langle M.P,M.Q,M.R\rangle$ after acting $M$ on the theta null-point. Here, to state a necessary and sufficient condition that $M.G=G$, we define as follows:
Definition 3.4.
For an integer $n\geq 1$, the set
$\bigl{\{}M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}\in{\rm Sp}_{6}(\mathbb{Z})\mid\gamma\equiv\mathbf{0}_{3}\!\pmod{n}\bigr{\}}\vspace{0.5mm}$
is a subgroup of ${\rm Sp}_{6}(\mathbb{Z})$, and we denote it by $\varGamma_{0}(n)$.
[…]
Remark 3.5.
There exist exactly $N_{3}(2)=135$ different $(2,2,2)$-isogenies from $A$ by Proposition 2.4, and thus the order of ${\rm Sp}_{6}(\mathbb{Z})/\varGamma_{0}(2)$ is equal to $135$.
For the reader’s convinience, we give a concrete algorithm for generating the list consisting of all different elements of ${\rm Sp}_{6}(\mathbb{Z})/\varGamma_{0}(2)$ as follows:
1:None
2:The list consisting of all different elements of ${\rm Sp}_{6}(\mathbb{Z})/\varGamma_{0}(2)$
3:Initialize $\mathcal{M}\leftarrow\emptyset$
4:repeat
5:     Generate a symplectic matrix $M\in{\rm Sp}_{6}(\mathbb{Z})$ randomly
6:     if $MN^{-1}\notin\varGamma_{0}(2)$ for all $N\in\mathcal{M}$ then
7:         Add $M$ to the list $\mathcal{M}$
8:     end if
9:until #$\#\mathcal{M}=135$
10:return $\mathcal{M}$
Algorithm 3.3  CollectSymplecticMatrices
Remark 3.6.
This list $\mathcal{M}$ does not depend on the characteristic $p$ of a field or an abelian threefold $A$. Hence, we only need to run Algorithm 3.3 once, and the result can be used many times thereafter.
In the following, let $\mathcal{M}$ be the output of the function CollectSymplecticMatrices. Then, we can compute theta null-points of $135$ abelian threefolds $(2,2,2)$-isogenous to $A$ by using Theorem 2.9.
1:A theta null-point $[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}$ of an abelian threefold $A$
2:A list of theta null-points of all abelian threefolds $(2,2,2)$-isogenous to $A$
3:Initialize $\mathcal{B}\leftarrow\emptyset$
4:for all $M\in\mathcal{M}$ do
5:     Write $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}$ where $\alpha,\beta,\gamma$ and $\delta$ are all $(3\times 3)$-matrices
6:     for $i\in I_{\rm even}$ do
7:         Let ${\vartheta^{\prime}}_{\!\!i}^{2}$ be the value determined from the action in Theorem 2.9
8:     end for
9:     Add ComputeOneRichelotIsogeny$([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$ to the list $\mathcal{B}$
10:end for
11:return $\mathcal{B}$
Algorithm 3.4  ComputeAllRichelotIsogenies
3.3. How to restore a defining equation of a genus-3 curve
In Section 2.3, we introduced some relations between a genus-3 curve $C$ and a theta null-point of the Jacobian of $C$. However, these are not enough to recover $C$ from a given theta null-point. Specifically, there are the following problems as it stands:
  • If $N_{\rm van}=1$ and $S_{\rm van}\neq\{61\}$, how should we restore a definind equation of the corresponding genus-3 curve?
  • If $N_{\rm van}=0$, we have to take some square roots of $\vartheta_{k}^{2}$ to use Theorem 2.17. How should we determine them?
In this subsection, we explain in detail how to do this.
Definition 3.7.
For each $i\in I_{\rm even}$, we write $\vartheta_{i}=\theta\bigl{[}\begin{smallmatrix}a_{1}&a_{2}&a_{3}\\ b_{1}&b_{2}&b_{3}\end{smallmatrix}\bigr{]}(0,\varOmega)$ and define
$P_{i}\coloneqq\begin{pmatrix}\mathbf{1}_{3}&\beta_{i}\\[-0.85358pt] \gamma_{i}&\mathbf{1}_{3}\end{pmatrix},\vspace{-1mm}$
where
  • If $i\notin\{27,31,45,47,54,55,59,61,62\}$, then
    $\beta_{i}\coloneqq{\rm diag}(b_{1},b_{2},b_{3}),\quad\gamma_{i}\coloneqq{\rm diag}(a_{1},a_{2},a_{3}).\vspace{0.5mm}$
  • Otherwise, we define as follows:
    $\beta_{27}\coloneqq{\small\begin{pmatrix}1&1&0\\ 1&1&0\\ 0&0&0\end{pmatrix}},\quad\gamma_{27}\coloneqq{\small\begin{pmatrix}1&-1&0\\ -1&1&0\\ 0&0&0\end{pmatrix}}.$
    $\beta_{31}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}},\quad\gamma_{31}\coloneqq{\small\begin{pmatrix}1&-1&0\\ -1&1&0\\ 0&0&0\end{pmatrix}}.$
    $\beta_{45}\coloneqq{\small\begin{pmatrix}1&0&1\\ 0&0&0\\ 1&0&1\end{pmatrix}},\quad\gamma_{45}\coloneqq{\small\begin{pmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{pmatrix}}.$
    $\beta_{47}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}},\quad\gamma_{47}\coloneqq{\small\begin{pmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{pmatrix}}.$
    $\beta_{54}\coloneqq{\small\begin{pmatrix}0&0&0\\ 0&1&1\\ 0&1&1\end{pmatrix}},\quad\gamma_{54}\coloneqq{\small\begin{pmatrix}0&0&0\\ 0&1&-1\\ 0&-1&1\end{pmatrix}}.$
    $\beta_{55}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}},\quad\gamma_{55}\coloneqq{\small\begin{pmatrix}0&0&0\\ 0&1&-1\\ 0&-1&1\end{pmatrix}}.$
    ${\small\beta_{59}\coloneqq\begin{pmatrix}1&-1&0\\ -1&1&0\\ 0&0&0\end{pmatrix}},\quad\gamma_{59}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}}.$
    $\beta_{61}\coloneqq{\small\begin{pmatrix}1&0&-1\\ 0&0&0\\ -1&0&1\end{pmatrix}},\quad\gamma_{61}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}}.$
    $\beta_{62}\coloneqq{\small\begin{pmatrix}0&0&0\\ 0&1&-1\\ 0&-1&1\end{pmatrix}},\quad\gamma_{62}\coloneqq{\small\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}}.$
Lemma 3.8.
For each $i\in I_{\rm even}$, the matrix $P_{i}$ defined as above belongs to ${\rm Sp}_{6}(\mathbb{Z})$. In addition, if we suppose that $\vartheta_{0}=0$, then we obtain $\vartheta_{i}(0,P_{i}\hskip 0.85358pt.\hskip 0.28453pt\varOmega)=0$.
Proof.
One can check that $\beta_{i},\gamma_{i}$ are both symmetric matrices and $\beta_{i}\gamma_{i}=\gamma_{i}\beta_{i}=\mathbf{0}_{3}$, which means that $P_{i}\in{\rm Sp}_{6}(\mathbb{Z})$ by Lemma 2.2. In addition for $a=b=0$, we have
$P_{i}\hskip 0.85358pt.\hskip 0.56905pta$ $=a\hskip 0.85358pt^{t}\delta_{i}-b\hskip 1.13809pt^{t}\gamma_{i}+\frac{1}{2}(\gamma_{i}\hskip 0.28453pt^{t}\delta_{i})_{0}=\frac{1}{2}(\gamma_{i}\hskip 0.28453pt^{t}\delta_{i})_{0}=(a_{1},a_{2},a_{3}),$
$P_{i}\hskip 0.85358pt.\hskip 0.56905ptb$ $=b\hskip 1.42262pt^{t}\alpha_{i}-a\hskip 0.85358pt^{t}\beta_{i}+\frac{1}{2}(\alpha_{i}\hskip 0.28453pt^{t}\beta_{i})_{0}=\frac{1}{2}(\alpha_{i}\hskip 0.28453pt^{t}\beta_{i})_{0}=(b_{1},b_{2},b_{3})$
by the equation (2.1). Then, this lemma follows from Theorem 2.9. ∎
In the following, let $A$ be an abelian threefold isomorphic to the Jacobian of a genus-3 hyperelliptic curve $C$, and we assume that a theta null-point of $A$ is given. In this setting, there uniquely exists $i\in I_{\rm even}$ such that $\vartheta_{i}=0$ by Proposition 2.15. If $i=61$, we can apply Proposition 2.16 for restoring a defining equation of $C$, but otherwise not. Here, let $M:=P_{61}{P_{i}}^{-1}\in{\rm Sp}_{6}(\mathbb{Z})$ and $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ be the theta null-point after acting $M$ on given $[\vartheta_{i}^{2}]_{i}$, and we have ${\vartheta^{\prime}}_{\!\!61}=0$ by Lemma 3.8. Then we can apply Proposition 2.16 to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$, which enables us to recover $C$ (Algorthm 3.5).
1:A theta null-point $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ with $N_{\rm van}=1$ of an abelian threefold $A$
2:A hyperelliptic curve $C$ of genus $3$ such that $A\cong{\rm Jac}(C)$
3:Let $i\in I_{\rm even}$ such that $\vartheta_{i}^{2}=0$ and define $M:=P_{61}{P_{i}}^{-1}$
4:Write $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}$ where $\alpha,\beta,\gamma$ and $\delta$ are all $(3\times 3)$-matrices
5:for $i\in I_{\rm even}$ do
6:     Let ${\vartheta^{\prime}}_{\!\!i}^{2}$ be the value determined from the action in Theorem 2.9
7:end for
8:Compute $\lambda_{1},\ldots,\lambda_{5}$ from the theta null-point $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}\in\mathbb{P}^{63}$ by the equation (2.2)
9:return the curve $y^{2}=x(x-1)\prod_{k=1}^{5}(x-\lambda_{k})$
Algorithm 3.5  RestoreHyperellipticCurve
[…]
(3.7) $\left\{\begin{array}[]{l}\vartheta_{0}\vartheta_{16}\vartheta_{45}\vartheta_{61}=\vartheta_{5}\vartheta_{21}\vartheta_{40}\vartheta_{56}-\vartheta_{12}\vartheta_{28}\vartheta_{33}\vartheta_{49},\\ \vartheta_{12}\vartheta_{21}\vartheta_{33}\vartheta_{56}=\vartheta_{5}\vartheta_{28}\vartheta_{40}\vartheta_{49}-\vartheta_{2}\vartheta_{27}\vartheta_{47}\vartheta_{54},\\ \vartheta_{3}\vartheta_{20}\vartheta_{32}\vartheta_{55}=\vartheta_{2}\vartheta_{21}\vartheta_{33}\vartheta_{54}+\vartheta_{12}\vartheta_{27}\vartheta_{47}\vartheta_{56},\end{array}\right.$
(3.8) $\left\{\begin{array}[]{l}\vartheta_{7}\vartheta_{14}\vartheta_{35}\vartheta_{42}=\vartheta_{5}\vartheta_{12}\vartheta_{33}\vartheta_{40}-\vartheta_{21}\vartheta_{28}\vartheta_{49}\vartheta_{56},\\ \vartheta_{14}\vartheta_{16}\vartheta_{35}\vartheta_{61}=\vartheta_{2}\vartheta_{28}\vartheta_{47}\vartheta_{49}-\vartheta_{5}\vartheta_{27}\vartheta_{40}\vartheta_{54},\\ \vartheta_{7}\vartheta_{16}\vartheta_{42}\vartheta_{61}=\vartheta_{2}\vartheta_{21}\vartheta_{47}\vartheta_{56}-\vartheta_{12}\vartheta_{27}\vartheta_{33}\vartheta_{54}\end{array}\right.$
for appropriate inputs $m_{1},m_{2},m_{3},m_{4}$. Squaring both sides of (3.7), we obtain
(3.9) $\begin{split}2\vartheta_{5}\vartheta_{12}\vartheta_{21}\vartheta_{28}\vartheta_{33}\vartheta_{40}\vartheta_{49}\vartheta_{56}&=\vartheta_{5}^{2}\vartheta_{21}^{2}\vartheta_{40}^{2}\vartheta_{56}^{2}+\vartheta_{12}^{2}\vartheta_{28}^{2}\vartheta_{33}^{2}\vartheta_{49}^{2}-\vartheta_{0}^{2}\vartheta_{16}^{2}\vartheta_{45}^{2}\vartheta_{61}^{2},\\ 2\vartheta_{2}\vartheta_{5}\vartheta_{27}\vartheta_{28}\vartheta_{40}\vartheta_{47}\vartheta_{49}\vartheta_{54}&=\vartheta_{5}^{2}\vartheta_{28}^{2}\vartheta_{40}^{2}\vartheta_{49}^{2}+\vartheta_{2}^{2}\vartheta_{27}^{2}\vartheta_{47}^{2}\vartheta_{54}^{2}-\vartheta_{12}^{2}\vartheta_{21}^{2}\vartheta_{33}^{2}\vartheta_{56}^{2},\\ 2\vartheta_{2}\vartheta_{12}\vartheta_{21}\vartheta_{27}\vartheta_{33}\vartheta_{47}\vartheta_{54}\vartheta_{56}&=\vartheta_{3}^{2}\vartheta_{20}^{2}\vartheta_{32}^{2}\vartheta_{55}^{2}-\vartheta_{2}^{2}\vartheta_{21}^{2}\vartheta_{33}^{2}\vartheta_{54}^{2}-\vartheta_{12}^{2}\vartheta_{27}^{2}\vartheta_{47}^{2}\vartheta_{56}^{2}.\end{split}$
By using these equations, we can restore a defining equation of a plane quartic $C$, as in Algorithm 3.6.
1:A theta null-point $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ with $N_{\rm van}=0$ of an abelian threefold $A$
2:A plane quartic $C$ of genus $3$ such that $A\cong{\rm Jac}(C)$
3:Compute ${a_{11}}^{\!\!\!2},\,{a_{12}}^{\!\!\!2},\,{a_{13}}^{\!\!\!2}$ by the equation (2.3) and choose their square roots
4:Compute $\vartheta_{5}\vartheta_{12}\vartheta_{33}\vartheta_{40},\,\vartheta_{5}\vartheta_{27}\vartheta_{40}\vartheta_{54},\,\vartheta_{12}\vartheta_{27}\vartheta_{33}\vartheta_{54}$ from values $a_{11},a_{21},a_{31}$
5:Compute $\vartheta_{21}\vartheta_{28}\vartheta_{49}\vartheta_{56},\,\vartheta_{2}\vartheta_{28}\vartheta_{47}\vartheta_{49},\,\vartheta_{2}\vartheta_{12}\vartheta_{33}\vartheta_{56}$ from the equation (3.9)
6:Compute $\vartheta_{7}\vartheta_{14}\vartheta_{35}\vartheta_{42},\,\vartheta_{14}\vartheta_{16}\vartheta_{35}\vartheta_{61},\,\vartheta_{7}\vartheta_{16}\vartheta_{42}\vartheta_{61}$ from the equation (3.8)
7:Compute $a_{12},a_{13},a_{22},a_{23},a_{32},$ and $a_{33}$ from the equation (2.3)
8:Compute $k_{1},k_{2},k_{3}$ from the equation (2.4)
9:Compute homogenous linear polynomials $\xi_{1},\xi_{2},\xi_{3}$ from Theorem 2.17
10:return the curve $(\xi_{1}x+\xi_{2}y-\xi_{3}z)^{2}-4\xi_{1}\xi_{2}xy=0$
Algorithm 3.6  RestorePlaneQuartic
3.4. Algorithm for computing the graph $\mathcal{G}_{3}(2,p)$
In this subsection, let $p$ be a prime integer greater than $7$ and we give an explicit algorithm for computing the superspecial Richelot isogeny graph $\mathcal{G}_{3}(2,p)$. First, we need to prepare the list $\Lambda_{2,p}$ of all isomorphism classes of superspecial genus-2 curves, in addition to the list $\Lambda_{1,p}$ of all isomorphism classes of supersingular elliptic curves. Using these lists, we can compute the graph $\mathcal{G}_{3}(2,p)$ as in Algorithm 3.7.
1:A prime integer $p\geq 11$ and the lists $\Lambda_{1,p}$ and $\Lambda_{2,p}$
2:The lists representing all vertices and all nodes of the graph $\mathcal{G}_{3}(2,p)$
3:Initialize $\mathcal{V}_{1}\leftarrow\emptyset,\,\mathcal{V}_{2}\leftarrow\emptyset,\,\mathcal{V}_{3}\leftarrow\emptyset,\,\mathcal{V}_{4}\leftarrow\emptyset$ and $\mathcal{S}\leftarrow\emptyset$
4:for all $E_{1},E_{2},E_{3}\in\Lambda_{1,p}$ do
5:     Let $j_{1},j_{2},j_{3}$ be the $j$-invariant of $E_{1},E_{2},E_{3}$
6:     Compute a theta null-point $[\vartheta^{2}_{i}]_{i}$ of the abelian threefold $E_{1}\times E_{2}\times E_{3}$
7:     Add $inv\coloneqq(j_{1},j_{2},j_{3})$ to the list $\mathcal{V}_{4}$ and $\bigl{(}[\vartheta^{2}_{i}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the list $\mathcal{S}$
8:end for
9:for all $E\in\Lambda_{1,p},\hskip 0.85358ptC\in\Lambda_{2,p}$ do
10:     Let $j$ be the $j$-invariant of $E$, and $j^{\prime}$ be the Igusa invariant of $C$
11:     Compute a theta null-point $[\vartheta^{2}_{i}]_{i}$ of the abelian threefold $E\times{\rm Jac}(C)$
12:     Add $inv\coloneqq(j,j^{\prime})$ to the list $\mathcal{V}_{3}$ and $\bigl{(}[\vartheta^{2}_{i}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the list $\mathcal{S}$
13:end for
14:Initialize $\mathcal{E}\leftarrow\emptyset$ and let $k\leftarrow 1$
15:repeat
16:     Let $\bigl{(}[\vartheta^{2}_{i}]_{i},inv\bigr{)}$ be the $k$-th element of $\mathcal{S}$
17:     for all $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}\in\hskip 2.27621pt$ComputeAllRichelotIsogenies($[\vartheta^{2}_{i}]_{i}$do
18:         Compute the number $N_{\rm van}$ of indices $i\in I_{\rm even}$ such that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$
19:         if $N_{\rm van}=0$ then
20:              Let $inv^{\prime}$ be the Dixmier-Ohno invariant of ${\sf RestorePlaneQuartic}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
21:              if $inv^{\prime}\notin\mathcal{V}_{1}$ then
22:                  Add $inv^{\prime}$ to the list $\mathcal{V}_{1}$ and $\bigl{(}[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the end of $\mathcal{S}$
23:              end if
24:         else if $N_{\rm van}=1$ then
25:              Let $inv^{\prime}$ be the Shioda invariant of ${\sf RestoreHyperellipticCurve}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
26:              if $inv^{\prime}\notin\mathcal{V}_{2}$ then
27:                  Add $inv^{\prime}$ to the list $\mathcal{V}_{2}$ and $\bigl{(}[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i},\hskip 0.56905ptinv\bigr{)}$ to the end of $\mathcal{S}$
28:              end if
29:         else if $N_{\rm van}=6$ then
30:              Restore $E\in\Lambda_{1,p}$ and $C\in\Lambda_{2,p}$, s.t. $E\times{\rm Jac}(C)$ corresponds to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$
31:              Let $inv^{\prime}:=(j,j^{\prime})$ where $j(E)=j$ and the Igusa invariant of $C$ is $j^{\prime}$
32:         else
33:              Restore $E_{1},E_{2},E_{3}\in\Lambda_{1,p}$, s.t. $E_{1}\times E_{2}\times E_{3}$ corresponds to $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$
34:              Let $inv^{\prime}:=\{j_{1},j_{2},j_{3}\}$ where $j(E_{1})=j_{1},\,j(E_{2})=j_{2}$ and $j(E_{3})=j_{3}$
35:         end if
36:         Add $(inv,inv^{\prime})$ to the list $\mathcal{E}$
37:     end for
38:     Set $k\leftarrow k+1$
39:until $k\hskip 0.56905pt>\#\mathcal{S}$
40:return $\mathcal{V}_{1}\sqcup\mathcal{V}_{2}\sqcup\mathcal{V}_{3}\sqcup\mathcal{V}_{4}$ and $\mathcal{E}$
Algorithm 3.7  ComputeSspRichelotIsogenyGraph
Recall from Section 2.3 that an abelian threefold is classified as Type $1,2,3$ or $4$ and $\mathcal{V}_{i}$ in the above algorithm represent the list of vertices of each type, respctively. In Steps $4$ and $9$, we have to compute a theta null-point of an abelian threefold of Type $3$ and $4$. Here, in the following, we explain how to obtain them briefly:
Lemma 3.9.
Let $E=\mathbb{C}/(\mathbb{Z}+\tau\mathbb{Z})$ with $\tau\in\mathbb{C}$ be an elliptic curve over $\mathbb{C}$ defined by the equation $y^{2}=x(x-1)(x-t)$. Then, there exists $c\in\mathbb{C}^{\times}$ such that
$\theta\bigl{[}\begin{smallmatrix}0\\ 0\end{smallmatrix}\bigr{]}(0,\tau)^{2}=c,\quad\theta\bigl{[}\begin{smallmatrix}0\\ 1/2\end{smallmatrix}\bigr{]}(0,\tau)^{2}=c\sqrt{t},\quad\theta\bigl{[}\begin{smallmatrix}1/2\\ 0\end{smallmatrix}\bigr{]}(0,\tau)^{2}=c\sqrt{1-t}$
and moreover $\theta\bigl{[}\begin{smallmatrix}1/2\\ 1/2\end{smallmatrix}\bigr{]}(0,\tau)=0$.
Proof.
Apply Thomae’s formula (cf. [32, Theorem 8.1]) for $g=1$. ∎
[…]
Remark 3.10.
[…]
In Steps 15–35, we compute each of the $135$ edges starting from a vertex in turn. If the codomain corresponds to a hyperelliptic curve of genus $3$, then we compute its Shioda invariant (cf. [37]). If the codomain corresponds to a plane quartic, we compute its Dixmier-Ohno invariant (cf. [12] and [33]). These invariants are given by the polynomials whose degrees are bounded by a constant, and hence they can be computed in constant time.
Remark 3.11.
For plane quartics $C,C^{\prime}$ over a field of characteristic 00, it is shown that $C\cong C^{\prime}$ if and only if their Diximier-Ohno invariants are equal to each other. For ones over a field of positive characteristic this is not proven, but it is likely to work also in characteristic $p>7$, see [29] for details.
If $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ is corresonding to an abelian threefold $A$ of Type $3$, we need to recover an elliptic curve $E$ and a genus-2 curve $C$ such that $A\cong E\times{\rm Jac}(C)$. Let $S_{\rm van}$ be the set of indices $i\in I_{\rm even}$ such that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$, then it follows from Proposition 2.15 that we have that #$\#S_{\rm van}=6$. If $S_{\rm van}=\{27,31,45,47,59,61\}$, then we can restore theta null-points of $E$ and ${\rm Jac}(C)$ by using Lemma 2.8. Even if not, this can be done by acting an appropriate $M\in{\rm Sp}_{6}(\mathbb{Z})$ on $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ to satisfy $S_{\rm van}=\{27,31,45,47,59,61\}$, and thus we can do computations in Step 28. Also in the case that $A$ is of Type $4$, acting $M\in{\rm Sp}_{6}(\mathbb{Z})$ so that $S_{\rm van}=\{27,31,45,47,54,55,59,61,62\}$ satisfies, we can do computations in Step 31.
Remark 3.12.
In Step 34, we add $(inv,inv^{\prime})$ to the list $\mathcal{E}$ if there exists an edge from the vertex with the invariant $inv$ to the vertex with the invriant $inv^{\prime}$. If we want to count edges with multiplicities, all we need to do is change $\mathcal{E}$ to the list with mulitplicity.
Summing up the above discussions, we obtain Algorithm 3.7. We remark that our algorithm is still valid over a field of characteristic $p\geq 11$. In the following, we assume the following assertions:
Assumption 3.13.
Let $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ be a theta null-points of a superspecial abelian threefold. Then, we can take each $\vartheta_{i}$ such that $\vartheta_{i}\in\mathbb{F}_{p^{2}}$.
We have confirmed that this assertion is correct for all $p<100$. Moreover, we have also experimented with some superspecial abelian threefolds for $p<10000$, and no counterexamples have been found to date.
Proof of Theorem 1.1.
Under the assumption stated above, we see that all computations in Algorithm 3.7 can be done over $\mathbb{F}_{p^{2}}$. It follows from
#$\#\Lambda_{1,p}=O(p),\quad\#\Lambda_{2,p}=O(p^{3})\quad{\rm and}\quad\#\Lambda_{3,p}=O(p^{6})$
that we can do Steps 2–6 within the complexity $\widetilde{O}(p^{3})$, and Steps 7–11 within the complexity $\widetilde{O}(p^{4})$. Next, the loops in Steps 13–37 are done $O(p^{6})$ times. For each loop, Steps 15–35 are executed exactly $135$ times, and moreover all operations can be performed in polynomial time of $\log{p}$. Therefore, Algorithm 3.7 can be done with $\widetilde{O}(p^{6})$ operations in the finite field $\mathbb{F}_{p^{2}}$. ∎
4. Computational results
In this section, we give some computational results obtained by executing main algorithm. We implemented the algorithms with Magma V2.26-10 and ran it on a machine with an AMD EPYC $7742$ CPU and $2$ TB of RAM.
4.1. Enumeration of superspecial hyperelliptic genus-3 curves
In this subsection, we enumerate superspecial hyperelliptic genus-3 curves in characteristic $p$ with $10011\leq p<$ 100. Indeed, we can do this by modifying Algorithm 3.7 to output the numbers of elements of $\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}$, and $\mathcal{V}_{4}$. Here, remark that #$\#\mathcal{V}_{1}$ (resp. #$\#\mathcal{V}_{2}$) is equal to the number of isomorphism classes of superspecial non-hyperelliptic (resp. hyperelliptic) curves of genus $3$.
Remark 4.1.
As there is no need to output the list $\mathcal{E}$, and therefore Steps $16,18$, and $20$ can be ignored. Moreover, the condition $k\hskip 0.56905pt>\#\mathcal{S}$ in Step $23$ can be replaced by the condition #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$, where #$\#\Lambda_{3,p}$ is the number of superspecial genus-3 curves (this number is obtained from [3, Theorem 3.10]). However, for the purpose of verifying the validity of our algorithm, we executed it without changing the condition $k\hskip 0.56905pt>\#\mathcal{S}$, in the following experiment.
The following table shows the cardinalities of $\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3},\mathcal{V}_{4}$ and the time it took. Here, the column total (resp. compl) denotes the number of times Step 13–37 are performed until the condition $k\hskip 0.56905pt>\#\mathcal{S}$ (resp. #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$) is satisfied. We note that ${\tt total}$ is equal to the sum of #$\#\mathcal{V}_{1},\#\mathcal{V}_{2},\#\mathcal{V}_{3}$, and #$\#\mathcal{V}_{4}$.
char.   #$\#\mathcal{V}_{1}$   #$\#\mathcal{V}_{2}$   #$\#\mathcal{V}_{3}$   #$\#\mathcal{V}_{4}$ total compl Time (sec)
$p=11$ 10 1 4 4 19 7 9.660
$p=13$ 18 1 3 1 23 5 7.790
$p=17$ 54 2 10 4 70 24 19.320
$p=19$ 87 4 14 4 109 48 29.860
$p=23$ 213 9 30 10 262 70 66.070
$p=29$ 681 10 54 10 755 249 189.090
$p=31$ 950 27 60 10 1047 249 258.710
$p=37$ 2448 35 93 10 2586 591 615.930
$p=41$ 4292 54 160 20 4526 1754 1070.190
$p=43$ 5567 82 180 20 5849 2685 1385.360
$p=47$ 9138 125 285 35 9583 3277 2295.400
$p=53$ 18032 153 390 35 18610 6663 4484.100
$p=59$ 33204 299 624 56 34183 11964 8719.060
$p=61$ 40259 262 565 35 41121 13015 10670.860
$p=67$ 69132 451 870 56 70509 25780 19646.600
$p=71$ 96717 647 1190 84 98638 30861 29135.890
$p=73$ 113778 582 1098 56 115514 35823 34888.620
$p=79$ 180273 942 1596 84 182895 51592 59625.370
$p=83$ 240755 1136 2080 120 244091 70445 83296.450
$p=89$ 362720 1402 2528 120 366770 122626 134065.830
$p=97$ 602062 2002 3200 120 607384 208415 230401.380
Table 1. The number of elements of lists $\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}$ and $\mathcal{V}_{4}$
Notably, the final results obtained show that #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$, and therefore our result is consistent with Brock’s result [3, Theorem 3.10].
Remark 4.2.
For the cardinalities of $\mathcal{V}_{3}$ and $\mathcal{V}_{4}$, we have
#$\#\mathcal{V}_{3}=\#\Lambda_{1,p}\cdot\#\Lambda_{2,p},\quad\#\mathcal{V}_{4}=\binom{\#\Lambda_{1,p}+2}{3}.$
Here, the numbers #$\#\Lambda_{1,p}$ and #$\#\Lambda_{2,p}$ are obtained by [11] and [21].
As can be seen in Table 1, the ratio ${\tt compl}/{\tt total}$ is approximately equal to $1/3$ for each $p$. Hence, if the termination condition is changed to #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$, and then the time required for computing #$\#\mathcal{V}_{2}$ is considered to be only $1/3$ of the time written in Table 1. In that case, listing up all superspecial genus-3 curves in characteristic $p=97$ takes about a day.
4.2. The existence of a superspecial hyperellitptic genus-3 curve
In this subsection, we check that there exists at least one superspecial hyperelliptic curve of genus $3$ in characteristic $p$ with $100007\leq p<$ 10000. First, the sufficient condition for the characteristic $p$ for such a curve to exist is given by the following proposition:
Theorem 4.3 ([3, Theorem 3.15] or [40, Theorem 2]).
  1. (1)
    The curve $y^{2}=x^{7}-1$ is superspecial if and only if $p\equiv 6\pmod{7}$.
  2. (2)
    The curve $y^{2}=x^{7}-x$ is superspecial if and only if $p\equiv 3\pmod{4}$.
  3. (3)
    The curve $y^{2}=x^{8}-1$ is superspecial if and only if $p\equiv 7\pmod{8}$.
By the above theorem, there exists a superspecial hyperelliptic curve of genus $3$ if $p\equiv 6\pmod{7}$ or $p\equiv 3\pmod{4}$. To investigate the other cases, we implemented the probabilistic algorithm (Algorithm 4.8) which outputs such a curve modifying Algorithm 3.7 slightly:
1:A prime integer $p\geq 7$
2:A superspecial hyperelliptic curve of genus $3$ in characteristic $p$
3:Generate a supersingular elliptic curve $E$ and $\mathcal{M}\leftarrow{\sf CollectSymplecticMatrices}()$
4:Compute a theta null-point $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ of $E^{3}$ by using Lemma 3.9 and Lemma 2.8
5:repeat
6:     Set $[\vartheta_{i}^{2}]_{i}\leftarrow[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ and choose a symplectic matrix $M\in\mathcal{M}$ randomly
7:     Write $M=\begin{pmatrix}\alpha&\beta\\[-0.85358pt] \gamma&\delta\end{pmatrix}$, where $\alpha,\beta,\gamma$, and $\delta$ are $(3\times 3)$-matrices
8:     for $i\in I_{\rm even}$ do
9:         Let ${\vartheta^{\prime}}_{\!\!i}^{2}$ be the value determined from the action in Theorem 2.9
10:     end for
11:     Compute the number $N_{\rm van}$ of indices $i\in I_{\rm even}$ such that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$
12:until $N_{\rm van}=1$
13:return ${\sf RestoreHyperellipticCurve}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
Algorithm 4.8  GenerateSingleSSpHypGenus3Curve
We can obtain the defining equation of a superspecial hyperelliptic genus-3 curve in the following procedure:
  • If $p\equiv 6\pmod{7}$, then output the curve $y^{2}=x^{7}-1$.
  • If $p\equiv 3\pmod{4}$, then output the curve $y^{2}=x^{7}-x$.
  • Otherwise, execute the Algorithm 4.8.
We remark that Algorithm 4.8 may return different outputs depending on randomness in Steps $1$ and $4$.
Executing the above procedure, we have checked that a superspecial hyperelliptic genus-3 curve exists for all characteristic $100007\leq p<$ 10000. As Algorithm 4.8 performs a random walk on $\mathcal{G}_{3}(2,p)$, the time to output such a curve varies, but we will list them below for reference:
  • For each $100007\leq p<$ 10000, the above procedure ends within 10,000 seconds. Moreover for almost $p$, it took less than 3,600 seconds.
  • It took 9,952 seconds for $p=8893$, which is the maximal time for $p<10000$.
  • It took very short time for some $p$ (e.g. about $30$ seconds for $p=9209$).
4.3. Drawing the graph $\mathcal{G}_{3}(2,p)$ for small $p$
As a direct result obtained from executing main algorithm, we draw the graph $\mathcal{G}_{3}(2,p)$ for $11\leq p\leq 19$ as follows:
Figure 1. The graph $\mathcal{G}_{3}(2,11)$
Figure 1. The graph $\mathcal{G}_{3}(2,11)$
Figure 2. The graph $\mathcal{G}_{3}(2,13)$
Figure 3. The graph $\mathcal{G}_{3}(2,17)$
Figure 3. The graph $\mathcal{G}_{3}(2,17)$
Figure 4. The graph $\mathcal{G}_{3}(2,19)$
In the above figures, we write abelian threefolds of Types $1$, $2$, $3$, and $4$ are written as blue, green, yellow, and red vertices respectively. Moreover, to improve visibility, we draw all the graphs as undirected graphs.
5. Concluding remarks
We proposed an algorithm (Algorithm 3.7) for computing the Richelot isogeny graph of superspecial abelian threefolds in characteristic $p>7$. As applications of this algorithm, we succeeded in listing up superspecial genus-3 curves for $p<100$ (Section 4.1), and finding one superspecial hyperelliptic curve of genus 3 for all $p$ with $100007\leq p<$ 10000 (Section 4.2). Our implementation with the Magma algebraic system [2] for listing up superspecial genus-3 curves is available at
https://github.com/Ryo-Ohashi/sspg3list.
The complexity of our algorithm is estimated as $\widetilde{O}(p^{6})$ operations in the field $\mathbb{F}_{p^{2}}$ under Remark 3.11 and Assumption 3.13, but these are open problems. Therefore, theoretical proofs of them are future works.
As other future works, it would be interesting to list up superspecial curves of genus $3$ in characteristic $p>100$. However, since enumerating in $p=97$ took about a day, enumerating even in $p=199$ can be expected to take more than two months. Hence, improving the algorithm in Section 4.1 is an important task. This requires an efficient search to obtain new nodes on the graph $\mathcal{G}_{3}(2,p)$, and then we need to investigate the structure of $\mathcal{G}_{3}(2,p)$ in detail.
On the other hand, we can expect from the results in Table 1 that
Expectation 5.1.
The number of isomorphism classes of superspecial hyperelliptic curves of genus $3$ in characteristic $p$ is $O(p^{4})$.
Under this expectation, the complexity of Algorithm 4.8 can be estimated as $\widetilde{O}(p^{2})$ operations in the field $\mathbb{F}_{p^{2}}$. A theoretical proof of the above expectation is also one of our future works.
Ryo Ohashi: Graduate School of Information Science and Technology, The University of Tokyo — 7-3-1 Hongo, Bunkyo-ku, Tokyo, 113-0033, Japan.
Hiroshi Onuki: Graduate School of Information Science and Technology, The University of Tokyo — 7-3-1 Hongo, Bunkyo-ku, Tokyo, 113-0033, Japan.
Momonari Kudo: Fukuoka Institute of Technology — 3-30-1 Wajiro-higashi, Higashi-ku, Fukuoka, 811-0295, Japan.
E-mail address: m-kudo@fit.ac.jp.
Ryo Yoshizumi: Joint Graduate Program of Mathematics for Innovation, Kyushu University — 744 Motooka, Nishi-ku, Fukuoka 819-0395, Japan.
Koji Nuida: Institute of Mathematics for Industry, Kyushu university — 744 Motooka, Nishi-ku, Fukuoka 819-0395, Japan / National Institute of Advanced Industrial Science and Technology (AIST) — 2-3-26 Aomi, Koto-ku, Tokyo, 135-0064, Japan.
E-mail address: nuida@imi.kyushu-u.ac.jp.